OpenStudy (anonymous):

Peyton has a credit card with an annual rate of 24.7% compounded monthly. She used the credit card to purchase cleaning supplies in the amount of $189.56. She can pay up to $72 on the credit card each month. How much total interest will she pay? $3.90 $7.47 $8.29

OpenStudy (tkhunny):

This is just not a good question. "up to" That is not meaningful. If she pays $10 / month, it will be more interest. If she pays $70 / month, it will be less interest. How much WILL she pay?

OpenStudy (anonymous):

Erm, I'm not sure

OpenStudy (anonymous):

I'm so lost D:

OpenStudy (tkhunny):

Well, as is so often the case, lacking a good problem statement, we have to make a good assumption and solve whatever problem we manage to define. It is kind of reasonable to suggest that if she CAN pay $72 / month, she WILL pay $72 / month. The problem statement DOES NOT say this. Let's assume it and move on. i = 0.247 -- Annual Interest Rate to be applied monthly. j = i/12 = 0.0205833 -- Monthly Interest Rate r = (1+j) -- Monthly accumulation factor. Can you accept this one assumption ($72/month) and these three definitions (i, j, r)? I haven't tried to solve the problem, yet. I'm just building the tools!

OpenStudy (anonymous):

Yep sounds resonable

OpenStudy (tkhunny):

The we're ready to roll... 189.56r - 72 = 121.46 Do you believe the state of affairs after one month, having accumulated interest and having made one payment?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

You do month 2. Accrue more interest and make another payment.

OpenStudy (anonymous):

What do you mean?

OpenStudy (tkhunny):

You do for month #2 exactly what I did for month #1. Start with 121.46 and see where it leads.

OpenStudy (anonymous):

What did you do for month 1?

OpenStudy (tkhunny):

You said you believed it. 189.56r - 72 = 121.46

OpenStudy (anonymous):

I didn't know the statement related to the problem

OpenStudy (anonymous):

I'm not sure what to do

OpenStudy (tkhunny):

Start with the beginning balance 189.56 Accrue the interest for one month 189.56r Make a payment 189.56r - 72 Result is the remaining balance after the first month and the first payment. 121.46

OpenStudy (anonymous):

Okay, so I don't know what to do for month 2?

OpenStudy (tkhunny):

EXACTLY the same thing.

OpenStudy (anonymous):

Are you saying I plug in 24.7% for r?

OpenStudy (tkhunny):

No, we defined r up above. That's the deal with these problems. Simply keeping track of everything is the hardest part. r = 1 + 0.247/12 = 1.020583 It is a monthly accumulation factor. It can be used to calculate the total interest and basis at the end of one month. Note: I NEVER would say "plug in". There is a principle called "substitution" that would be applicable.

OpenStudy (anonymous):

Okay so 2 months would equal 2.002058333

OpenStudy (anonymous):

Is there any way we can finish this faster, I wantto go to bed in like 10 minutes and I need to wake up early, this is the last question I have to do x.x

OpenStudy (tkhunny):

No. Start with the beginning balance 121.46 Accrue the interest for one month 121.46r Make a payment 121.46r - 72 Result is the remaining balance after the second month and the second payment. You tell me what it is...

OpenStudy (anonymous):

Is the answer $3.90 ?

OpenStudy (anonymous):

At first I got $3.41 but that was wrong

OpenStudy (tkhunny):

You need to understand the idea of accruing interest and making periodic payments, You will have to tell me how you managed those answers. I do not care about guessing. Show your work and I can see what it is you are doing.

OpenStudy (anonymous):

I'd love to but unfortunately I don't have the time now, so I'll go with $3.90, thank you anyways

OpenStudy (tkhunny):

That's about half the correct answer. When you have the time to learn, feel free to come back.

OpenStudy (tkhunny):

Keep in mind that it is a horrible problem statement.

OpenStudy (anonymous):

Okay I'll come back tomorrow and explain how I got $3.41